# Impact of the Magnetic Horizon on the Interpretation of the Pierre Auger Observatory Spectrum and Composition Data

### v1

@@ Abstract
The flux of ultra-high energy cosmic rays reaching Earth above the ankle energy (5 EeV) can be described as a mixture of nuclei injected by extragalactic sources with very hard spectra and a low rigidity cutoff. Extragalactic magnetic fields existing between the Earth and the closest sources can affect the observed CR spectrum by reducing the flux of low-rigidity particles reaching Earth. We perform a combined fit of the spectrum and distributions of depth of shower maximum measured with the Pierre Auger Observatory including the effect of this magnetic horizon in the propagation of UHECRs in the intergalactic space. We find that, within a specific range of the various experimental and phenomenological systematics, the magnetic horizon effect can be relevant for turbulent magnetic field strengths in the local neighbourhood of order $B_{\rm rms}\simeq(50-100)\,{\rm nG}\,(20\rm{Mpc}/{d_{\rm s})(100\,\rm{kpc}/L_{\rm coh}})^{1/2}$, with $d_{\rm s}$ the typical intersource separation and $L_{\rm coh}$ the magnetic field coherence length. When this is the case, the inferred slope of the source spectrum becomes softer and can be closer to the expectations of diffusive shock acceleration, i.e., $\propto E^{-2}$. An additional cosmic-ray population with higher source density and softer spectra, presumably also extragalactic and dominating the cosmic-ray flux at EeV energies, is also required to reproduce the overall spectrum and composition results for all energies down to 0.6 EeV.
@@ 2 The source populations
The general framework we consider in this study to reproduce the energy spectrum both below and above the ankle feature is that already explored in [2], consisting of two different source populations. One source population dominates the CR flux above a few EeV while a second one dominates at lower energies. For each of these populations the source spectra of the different mass components are considered to be described as power laws having a rigidity dependent cutoff which strongly suppresses the fluxes above an energy $ZR_{\rm cut}$. In this case the differential particle generation rate of each component of atomic number $Z$ and mass number $A$, per unit volume, energy and time, is
@@
with the parameter $\Delta$ determining the steepness of the cutoff shape, which turns out to have a significant impact on the fit. We will consider the three representative values $\Delta=1,2$ and 3.222The case $\Delta=1$ is very close to the broken exponential cutoff case considered in [2].
@@ 3 The magnetic horizon effect
The presence of extragalactic magnetic fields (EGMF), possibly enhanced in our neighbourhood inside the Local Supercluster region where the closest UHECR sources may lie, can be the dominant contributor to a magnetic horizon effect and may hence affect the observed cosmic-ray spectrum. We will model for simplicity the EGMF as being turbulent and isotropic, parameterized by its root-mean-square amplitude ($B_{\rm rms}$) and the coherence length ($L_{\rm coh}$), considering a Kolmogorov spectrum for the turbulence. A critical energy can be defined as that for which the effective Larmor radius associated to $B_{\rm rms}$ equals the coherence length, and for a particle of atomic number $Z$ it is given by $E_{\rm crit}\equiv|e|ZB_{\rm rms}L_{\rm coh}\equiv ZR_{\rm crit}$, with the critical rigidity
@@
where $x\equiv E/(ZR_{\rm crit})$ and $X_{\rm s}=d_{\rm s}/\sqrt{r_{H}L_{\rm coh}}$ is the normalized intersource distance, with $r_{H}=c/H_{0}$ the Hubble radius (in terms of the the speed of light $c$ and the present day Hubble constant $H_{0}\simeq 70$ km s-1 Mpc-1), so that
@@
The magnetic suppression factor $G$ is displayed in Fig. 1 as a function of $E/E_{\rm crit}$, for different values of $X_{\rm s}$ and for the NE (left) and SFR (right) evolutions, considering $\gamma=1$. The implied strong hardening of the spectrum at energies $E\lesssim E_{\rm crit}\equiv ZR_{\rm crit}$ is apparent. This effect is clearly rigidity dependent, given its magnetic origin. One finds that for increasing $X_{\rm s}$ the suppression appears for larger values of $E/E_{\rm crit}$. Hence, increasing $X_{\rm s}$ and simultaneously decreasing $E_{\rm crit}$ the different curves can lead to comparable suppressions as a function of the energy, what will lead to some approximate degeneracies between these parameters, although the shape of the suppression becomes steeper for increasing values of $X_{\rm s}$. For the same values of the parameters $X_{\rm s}$ and $E_{\rm crit}$ the suppressions are milder for the SFR case than for the NE one, given that due to the enhanced emission at high redshifts the particles have on average more time to arrive from their sources in the first case (note that for diffusing particles the redshift is a measure of the time travelled rather than of the source distance). We show the results both for primaries (solid lines), secondary protons (dashed lines) and intermediate secondary nuclei (dotted lines), where the primaries are those in which the detected nucleus is in the same mass group as the emitted one, considering the different mass groups as those with values of $A$ of 1 (H), 2–4 (He), 5–16 (N), 17–30 (Si) and 31-56 (Fe). The intermediate secondary nuclei are those in a lighter mass group than the primary injected particle. The milder suppression of the secondaries is also understood because their average redshift of production in photodisintegration processes is higher than the average production redshift of the primaries, given that the photodisintegrations processes get enhanced at higher redshifts. Also note that the suppression of the secondary protons is quite similar to the suppression of the secondary nuclei of the same rigidity.
Refer to captionRefer to caption
@@ 4.4.2 Results obtained including the magnetic horizon effect
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 4.5 Impact of systematic uncertainties on the energy and $X_{\rm max}$ calibrations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 5 Conclusions
On the other hand, magnetic fields are ubiquitous in the Universe, and although their strength is poorly known, they are expected to be enhanced in large scale structures, in particular in our neighbourhood. Given the discreteness of the source distribution, a magnetic horizon effect can significantly modify the observed spectrum because the CRs may not be able to reach us at low rigidities, making the observed spectrum harder than the source one. We hence explored in this paper the impact on the combined fit results when taking into account the magnetic horizon effect. We included this effect only for the high-energy component, given its assumed lower source density, considering that the associated effects in the denser low-energy component happen at energies below those considered in the analysis. In particular, if the intersource separation of the low-energy component were an order of magnitude smaller than that of the high- energy one, the magnetic suppression effects on it would be very mild above 1 EeV. Let us note that taking into account that the inferred emissivity per unit volume and time for the low and high energy populations are of comparable magnitude, one expects in that case that the typical individual luminosities of the much more abundant LE sources be fainter by at least three orders of magnitude with respect to those of the HE sources.
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption

### v2

@@ Abstract
The flux of ultra-high energy cosmic rays reaching Earth above the ankle energy (5 EeV) can be described as a mixture of nuclei injected by extragalactic sources with very hard spectra and a low rigidity cutoff. Extragalactic magnetic fields existing between the Earth and the closest sources can affect the observed CR spectrum by reducing the flux of low-rigidity particles reaching Earth. We perform a combined fit of the spectrum and distributions of depth of shower maximum measured with the Pierre Auger Observatory including the effect of this magnetic horizon in the propagation of UHECRs in the intergalactic space. We find that, within a specific range of the various experimental and phenomenological systematics, the magnetic horizon effect can be relevant for turbulent magnetic field strengths in the local neighbourhood in which the closest sources lie of order $B_{\rm rms}\simeq(50-100)\,{\rm nG}\,(20\,\rm{Mpc}/{d_{\rm s})(100\,\rm{kpc}/L_{\rm coh}})^{1/2}$, with $d_{\rm s}$ the typical intersource separation and $L_{\rm coh}$ the magnetic field coherence length. When this is the case, the inferred slope of the source spectrum becomes softer and can be closer to the expectations of diffusive shock acceleration, i.e., $\propto E^{-2}$. An additional cosmic-ray population with higher source density and softer spectra, presumably also extragalactic and dominating the cosmic-ray flux at EeV energies, is also required to reproduce the overall spectrum and composition results for all energies down to 0.6 EeV.
@@ 2 The source populations
The general framework we consider in this study to reproduce the energy spectrum both below and above the ankle feature is that already explored in [2], consisting of two different source populations. One source population dominates the CR flux above a few EeV while a second one dominates at lower energies. For each of these populations the source spectra of the different mass components are considered to be described as power laws having a rigidity dependent cutoff which strongly suppresses the fluxes above an energy $ZR_{\rm cut}$. This kind of cutoffs are expected when the acceleration has an electromagnetic origin, so that the maximum energy achieved is proportional to the charge of the particle, and when there are no strong interactions in the acceleration region.222 A different situation could be that in which the maximum energy would result from interaction effects, since for instance nuclei could be limited by photodisintegration processes which do not affect the protons. On the other hand, scenarios in which the sources have a distribution of cutoff rigidities have been discussed in [25]. In that case, the differential particle generation rate of each component of atomic number $Z$ and mass number $A$, per unit volume, energy and time, is
@@
with the parameter $\Delta$ determining the steepness of the cutoff shape, which turns out to have a significant impact on the fit. We will consider the three representative values $\Delta=1,2$ and 3.333The case $\Delta=1$ is very close to the broken exponential cutoff case considered in [2].
@@ 3 The magnetic horizon effect
The presence of extragalactic magnetic fields (EGMF) can be the dominant contributor to a magnetic horizon effect and may hence affect the observed cosmic-ray spectrum. In particular, EGMF are expected to be enhanced in our local neighbourhood where the closest UHECR sources should lie, such as within the Local Supercluster which has a typical radial extent of about 20 Mpc. Also note that the assumption of equipartition between the energy in thermal gas motion and in the magnetic field leads to an estimation of the magnetic field strength within filaments and sheets of the large scale structure of order $10^{2}$ nG [44]. We will model for simplicity the EGMF as being turbulent and isotropic, parameterized by its root-mean-square amplitude ($B_{\rm rms}$) and coherence length ($L_{\rm coh}$), considering a Kolmogorov spectrum for the turbulence. A critical energy can be defined as that for which the effective Larmor radius associated to $B_{\rm rms}$ equals the coherence length, and for a particle of atomic number $Z$ it is given by $E_{\rm crit}\equiv|e|ZB_{\rm rms}L_{\rm coh}\equiv ZR_{\rm crit}$, with the critical rigidity
@@
where $x\equiv E/(ZR_{\rm crit})$ and $X_{\rm s}=d_{\rm s}/\sqrt{r_{H}L_{\rm coh}}$ is the normalized intersource distance, with $r_{H}=c/H_{0}$ the Hubble radius (in terms of the speed of light $c$ and the present day Hubble constant $H_{0}\simeq 70$ km s-1 Mpc-1), so that
@@
If one considers magnetic field amplitudes in the range $1\,{\rm nG}\lesssim B_{\rm rms}\lesssim 200$ nG and coherence length such that $25\,{\rm kpc}\lesssim L_{\rm coh}\lesssim 1$ Mpc, one expects that $0.022\,{\rm EeV}\lesssim R_{\rm crit}\lesssim 180$ EeV. Moreover, if the intersource average distance is in the range $4\,{\rm Mpc}\lesssim d_{\rm s}\lesssim 40$ Mpc one should have that $0.05\lesssim X_{\rm s}\lesssim 4$. We will hence consider parameters $R_{\rm crit}$ and $X_{\rm s}$ within these ranges when performing the fits that include the magnetic horizon effects.
The magnetic suppression factor $G$ is displayed in Fig. 1 as a function of $E/E_{\rm crit}$, for different values of $X_{\rm s}$ and for the NE (left) and SFR (right) evolutions, considering $\gamma=1$. The implied strong hardening of the spectrum at low energies is apparent, with the suppression occurring in general in the regime of resonant diffusion, i.e. for energies below $E_{\rm crit}\equiv ZR_{\rm crit}$. This effect is clearly rigidity dependent, given its magnetic origin. One also finds that for increasing $X_{\rm s}$ the suppression appears for larger values of $E/E_{\rm crit}$. Hence, increasing $X_{\rm s}$ and simultaneously decreasing $E_{\rm crit}$ the different curves can lead to comparable suppressions as a function of the energy, what will lead to some approximate degeneracies between these parameters, although the shape of the suppression becomes steeper for increasing values of $X_{\rm s}$. For the same values of the parameters $X_{\rm s}$ and $E_{\rm crit}$ the suppressions are milder for the SFR case than for the NE one, given that due to the enhanced emission at high redshifts the particles have on average more time to arrive from their sources in the first case (note that for diffusing particles the redshift is a measure of the time travelled rather than of the source distance). We show the results both for primaries (solid lines), secondary protons (dashed lines) and intermediate secondary nuclei (dotted lines), where the primaries are those in which the detected nucleus is in the same mass group as the emitted one, considering the different mass groups as those with values of $A$ of 1 (H), 2–4 (He), 5–16 (N), 17–30 (Si) and 31-56 (Fe). The intermediate secondary nuclei are those in a lighter mass group than the primary injected particle. The milder suppression of the secondaries is also understood because their average redshift of production in photodisintegration processes is higher than the average production redshift of the primaries, given that the photodisintegrations processes get enhanced at higher redshifts. Also note that the suppression of the secondary protons is quite similar to the suppression of the secondary nuclei of the same rigidity.
@@ 5 Conclusions
On the other hand, magnetic fields are ubiquitous in the Universe, and although their strength is poorly known, they are expected to be enhanced in large scale structures, in particular in the Local Supercluster. Given the discreteness of the source distribution, a magnetic horizon effect can significantly modify the observed spectrum because the CRs may not be able to reach us at low rigidities, making the observed spectrum harder than the source one. We hence explored in this paper the impact on the combined fit results when taking into account the magnetic horizon effect. We included this effect only for the high-energy component, given its assumed lower source density, considering that the associated effects in the denser low-energy component happen at energies below those considered in the analysis. In particular, if the intersource separation of the low-energy component were an order of magnitude smaller than that of the high- energy one, the magnetic suppression effects on it would be very mild above 1 EeV. Let us note that taking into account that the inferred emissivity per unit volume and time for the low and high energy populations are of comparable magnitude, one expects in that case that the typical individual luminosities of the much more abundant LE sources be fainter by at least three orders of magnitude with respect to those of the HE sources.